# General-purpose post-sampling reweighting method for multimodal target measures

### v1

@@ 1 Introduction
In the present work, only point (iii) is addressed. This is based on the idea that, in many situations, it is in fact better to deal separately with the three parts of the problem. Due to dimensional complexity, solving point (i) in general is an extremely hard problem. In practice, each application has its model-dependent heuristics, exploiting additional structure or a-priori informations on the potential landscape. In any cases, when trying to solve (i) without dealing with (ii) and (iii), we do not have to keep track of the measure which is sampled by the method (contrary to e.g. importance sampling), so that we can use random initialisations with gradient descent (as is standard for non-convex optimization in machine learning), or completely non-equilibrium methods (e.g. saddle point search [26, 22]) or possibly adaptive methods (e.g. OPES [19], or Gaussian-accelerated simulations [7]) with very “aggressive” parameters (which would yield a poor sampling due to high weight variance). In some situations, the modes may even be known a priori (for instance in protein-ligand binding in molecular simulations [14]). In this paper we assume that step (i) has been addressed by some methods and that the modes are given. If some modes have been missed at this point, the method presented below will not find them. Moreover, we have in mind a situation where there is a no more than a few dozens of modes (corresponding e.g. to relevant macroscopic states in molecular simulations, each of which gathering possibly several close small-scale local energy minima).
As already mentioned, local MCMC are usually quite efficient for step (ii), so we assume now that we have samples in each mode. For instance, once the modes have been identified, we may simply run independent MCMC samplers initialized at each of these modes. Our goal is to perform step (iii) in this situation. Notice that, even without distinguishing the three steps, the question is also of interest in many cases with current practical methods (for instance adaptive importance sampling) because equilibration between modes happen at a time-scale larger than mode exploration and local sampling, which means that in practice the simulations often end with many modes being locally sampled but an incorrect relative weights between them.
@@
In summary, motivated by the above three-step viewpoint to multimodal sampling, our goal is to introduce a very simple general method for step (iii) when the modes have already been found and sampled. Contrary to most approaches we are aware of (for instance those reviewed in [17], or discussed in [24, Appendix A]), the method does not produce any new samples (or doesn’t add any new component to a mixture of parametrized distributions): this is a purely post-sampling step, which can be performed at the end of any sampling procedure. In other words, it addresses the following question: when our sampling budget is all spent, what is the best we can do with its output ?
@@
Since [10], a variety of methods have combined VI and sampling. Sometimes MCMC is used to provide a richer family of probability measures for VI (e.g. by propagating for a MCMC few steps a parametrized measure [32, 33], or by sampling latent variables from which the conditional densities of the observations are parametrized [18]) – our method can be seen like this, in an extreme way since the component of the mixture $\pi(p)$ are just directly given by sampling –, or VI is used to design (with mixture of Gaussian distributions, normalizing flows, etc.) good proposal kernels for importance sampling or MCMC [16, 39, 12], see also references within. In this last situation, and similarly in adaptive biasing importance sampling methods [19, 9], the idea is that, once the modes have been found, the adaptive mechanism will converge to a stationary Markovian dynamics (or and IS proposal for i.i.d. samples) which performs transitions between modes easily, either thanks to a proposal which allows for jumps between modes, or because some high-probability paths have been created between them. This means that, once this idealized sampler has been reached, we still have to wait for the Monte-Carlo estimation of the relative weights. By contrast, the point of view taken in the present work is that, since the local MCMC samples are good, we don’t expect to get more information by producing more samples and we estimate deterministically the weights.
@@ 3.2 Case with disjoint supports
(in this disjoint support case, we retrieve as discussed e.g. in [17, Section 2] the link between the mode weight estimation and the computation of the normalization constant $\mu(\mathcal{D}_{k})$ of the measure $\mathbbm{1}_{\mathcal{D}_{k}}\mu$). More generally the goal of the algorithm in this situation is to find the best approximation of $\mu_{\mathcal{D}}$.
@@
If $\nu_{k}=\mu_{k}$, the correct weights are recovered. However in general if $\nu_{k}\neq\mu_{k}$ there is still an error, which is expected since within a cluster the different samples are not re-weighted one relatively to the others. It is interesting to see that this asymptotic error is minimized by minimizing $\mathrm{KL}(\nu_{k}|\mu_{k})$ (which is in fact simply a consequence of the extensivity property of the relative entropy), which is precisely the quantity that the algorithms aims at minimizing at the global level. This means that, in multi-scale situations, it is consistent to apply the method hierarchically, i.e. combining small clusters together with re-weighting in small groups, and then merging some of these groups with a new relative re-weighting, etc. According to (14), taking the optimal weights at the smaller scale minimizes the error at the larger scales.
@@ 4 Numerical experiments
Fourth, to illustrate the capacity of the scheme to deal with more modes, we consider a mixture of $K=10$ Gaussian distributions, in dimension $d=100$. For the weights $p^{*}_{\mathrm{true}}$, we take $0.4$, $0.3$ and $0.1$ for the three first coordinates, and for the rest we take K3K-3 i.i.d. uniform random variables rescaled to sum up to $0.2$. For the parameters of each Gaussian distribution, half of them have variance $\Sigma_{1}$ and the other half has variance $\Sigma_{2}$, as given by (16). The means are generated as $K$ independent variable with law $\mathcal{N}(0,I_{d})$.

### v2

@@ 1 Introduction
In the present work, only point (iii) is addressed. This is based on the idea that, in many situations, it is in fact better to deal separately with the three parts of the problem. Due to dimensional complexity, solving point (i) in general is an extremely hard problem. In practice, each application has its model-dependent heuristics, exploiting additional structure or a-priori informations on the potential landscape. In any cases, when trying to solve (i) without dealing with (ii) and (iii), we do not have to keep track of the measure which is sampled by the method (contrary to e.g. importance sampling), so that we can use random initialisations with gradient descent (as is standard for non-convex optimization in machine learning), or completely non-equilibrium methods (e.g. saddle point search [29, 25]) or possibly adaptive methods (e.g. OPES [20], or Gaussian-accelerated simulations [8]) with very “aggressive” parameters (which would yield a poor sampling due to high weight variance). In some situations, the modes may even be known a priori (for instance in protein-ligand binding in molecular simulations [15]). In this paper we assume that step (i) has been addressed by some methods and that the modes are given. If some modes have been missed at this point, the method presented below will not find them. Moreover, we have in mind a situation where there is no more than a few dozens of modes (corresponding e.g. to relevant macroscopic states in molecular simulations, each of which gathering possibly several close small-scale local energy minima).
As already mentioned, local MCMC are usually quite efficient for step (ii), so we assume now that we have samples in each mode. For instance, once the modes have been identified, we may simply run independent MCMC samplers initialized at each of these modes. Our goal is to perform step (iii) in this situation. Notice that, even without distinguishing the three steps, the question is also of interest in many cases with current practical methods (for instance adaptive importance sampling) because equilibration between modes happen at a time-scale larger than mode exploration and local sampling, which means that in practice the simulations often end with many modes being locally sampled but an incorrect relative weights between them. In the vocabulary of statistical physics, step (iii) corresponds to estimating free energy differences. Standard methods such a Bennet acceptance ratio [23, 3] or targeted free energy [21, 45] require to have some overlap between the samples in each state or to have some transformation between them, which is not easily obtained when the states are very separated and no intermediate samples in the separating region are available.
@@
In summary, motivated by the above three-step viewpoint to multimodal sampling, our goal is to introduce a very simple general method for step (iii) when the modes have already been found and sampled. Contrary to most approaches we are aware of (for instance those reviewed in [18], or discussed in [27, Appendix A]), the method does not produce any new sample (or doesn’t add any new component to a mixture of parametrized distributions): this is a purely post-sampling step, which can be performed at the end of any sampling procedure. In other words, it addresses the following question: when our sampling budget is all spent, what is the best we can do with its output ?
@@
Since [11], a variety of methods have combined VI and sampling. Sometimes MCMC is used to provide a richer family of probability measures for VI (e.g. by propagating for a MCMC few steps a parametrized measure [36, 37], or by sampling latent variables from which the conditional densities of the observations are parametrized [19]) – our method can be seen like this, in an extreme way since the component of the mixture $\pi(p)$ are just directly given by sampling –, or VI is used to design (with mixture of Gaussian distributions, normalizing flows, etc.) good proposal kernels for importance sampling or MCMC [17, 46, 13], see also references within. In this last situation, and similarly in adaptive biasing importance sampling methods [20, 10], the idea is that, once the modes have been found, the adaptive mechanism will converge to a stationary Markovian dynamics (or an importance sampling proposal for i.i.d. samples) which performs transitions between modes easily, either thanks to a proposal which allows for jumps between modes, or because some high-probability paths have been created between them. This means that, once this idealized sampler has been reached, we still have to wait for the Monte-Carlo estimation of the relative weights. By contrast, the point of view taken in the present work is that, since the local MCMC samples are good, we don’t expect to get more information by producing more samples and we estimate deterministically the weights.
@@
In [34], densities are estimated in each mode, and then a mixture of these densities is used in a Metropolis-Hastings MCMC procedure. Contrary to our method, the weights of the mixture are adaptively estimated along the MCMC sampling.
@@ 3.2 Case with disjoint supports
(in this disjoint support case, we retrieve as discussed e.g. in [18, Section 2] the link between the mode weight estimation and the computation of the normalization constant $\mu(\mathcal{D}_{k})$ of the measure $\mathbbm{1}_{\mathcal{D}_{k}}\mu$). More generally the goal of the algorithm in this situation is to find the best approximation of $\mu_{\mathcal{D}}$.
@@
If $\nu_{k}=\mu_{k}$, the correct weights are recovered. However in general if $\nu_{k}\neq\mu_{k}$ there is still an error, which is expected since within a cluster the different samples are not re-weighted one relatively to the others. It is interesting to see that this asymptotic error is minimized by minimizing $\mathrm{KL}(\nu_{k}|\mu_{k})$ (which is in fact simply a consequence of the extensivity property of the relative entropy), which is precisely the quantity that the algorithm aims at minimizing at the global level. This means that, in multi-scale situations, it is consistent to apply the method hierarchically, i.e. combining small clusters together with re-weighting in small groups, and then merging some of these groups with a new relative re-weighting, etc. According to (14), taking the optimal weights at the smaller scale minimizes the error at the larger scales.
@@
When we use this dimension reduction, we do not need to reconstitute the variables $z$ from the variables $y$, contrary to methods which use the density estimation to design adaptive importance sampling or Metropolis-Hastings proposals [39, 38]. More generally, in our method in the non-overlapping case, the density estimation is an intermediate tool which is only used to estimate the entropy in each mode.
@@ 4 Numerical experiments
Fourth, to illustrate the capacity of the scheme to deal with more modes, we consider a mixture of $K=10$ Gaussian distributions, in dimension $d=100$. For the weights $p^{*}_{\mathrm{true}}$, we take $0.4$, $0.3$ and $0.1$ for the three first coordinates, and for the rest we take K3K-3 i.i.d. uniform random variables rescaled to sum up to $0.2$. For the parameters of each Gaussian distribution, half of them (clusters $k$ with odd $k$) have variance $\Sigma_{1}$ and the other half has variance $\Sigma_{2}$, as given by (16). The means are generated as $K$ independent variables with law $\mathcal{N}(0,I_{d})$.